# Do Large Language Models (Really) Need Statistical Foundations?

### v2

@@ 1 Introduction
Given that LLMs involve many statistical aspects, a burgeoning body of statistical research on LLMs has emerged (Ji et al., 2025). Nonetheless, the question we pose here is whether LLMs would genuinely benefit from statistical foundations that our community develops and could develop. Furthermore, we ask if such statistical contributions would lead to improved development, deployment, and application, particularly to guide and enhance their real-world use? In other words, we move beyond the discussion of how LLMs can be used to enhance statistical analysis and statistical education to focus on whether statistical methodology and insights can improve LLMs themselves. [^3] For clarity, this paper does not address how LLMs can be used to enhance statistical analysis or statistical education. Interested readers are referred to Tu et al. (2024).
@@ 2 LLMs as Statistical Models
While designed as a predictive method where the underlying data-generating mechanism is treated as unknown, LLMs exhibit characteristics that distinguish them significantly from many other predictive algorithms (Breiman, 2001; Efron, 2020). Predictive methods like decision trees or support vector machines often rely on feature engineering or, equivalently, carefully chosen kernels, whereas modern LLMs operate in a nearly end-to-end data-centric manner, trained directly on vast quantities of raw data with minimal human intervention beyond data selection and cleaning. [^4] Indeed, data curation is nontrivial in practice, but from a methodological perspective, no detailed human knowledge of linguistic constructs is explicitly hard-coded into LLMs at training time.
Owing to this strong data dependency, the capabilities of LLMs are largely determined by the properties and scale of the training data (Thrush et al., 2024). This fact is quantitatively captured by scaling laws, which demonstrate a predictable relationship between model capabilities and the volume of training data during the pre-training phase (Kaplan et al., 2020). In essence, scaling laws suggest that LLMs are what they are trained on, elevating data to the most critical component of their development. As evidence, while open-source LLM developers often release model weights and sometimes some technical details of their training processes, the composition and preparation of their training datasets are almost never disclosed (Grattafiori et al., 2024; Liu et al., 2024).
@@
The predominant training paradigm for modern LLMs is next-token prediction (Radford et al., 2018). [^5] While other objectives exist (e.g., masked language modeling in BERT) and new approaches are emerging (e.g., diffusion models for text), the core task remains generative. Crucially, this generative process is inherently stochastic, depending not only on the prior context but also on the data used for training LLMs. This randomness is not merely an artifact but arguably a necessity for modeling human language, which itself—whether in stories, manuals, essays, creative writing, code, or mathematical proofs—is generative and rarely follows a deterministic path. This contrasts sharply with many earlier deep learning applications, particularly in classification, where outputs often correspond to fixed, deterministic ground truths.
@@
Furthermore, the variability and uncertainty stemming from the stochastic nature of LLMs demand statistical analysis. Because LLM outputs are inherently variable, understanding and quantifying the uncertainty associated with their responses is critical. This is especially true in scientific applications where reproducibility is required, and in high-stakes decision-making, such as medical diagnostics, where the model’s confidence level can drastically influence subsequent actions. Statistics offers a rich toolkit for analyzing variability, uncertainty, and miscalibration, and subsequently reducing them.
@@ 3 LLMs as Black-box Models
The black-box nature is compounded by the fact that there is not one single “correct” architecture for LLMs achieving high performance. While the Transformer has been dominant, simplified Transformer and mixture-of-experts Transformer variants are shown empirically to work well (He and Hofmann, 2024; Dai et al., 2024). [^6] Notably, the Transformer can learn not only data patterns but also entire algorithms for statistical problems (He et al., 2025; Cao et al., 2025). Moreover, non-attention-based architectures like state-space models (Mamba) (Gu and Dao, 2024), recurrent structures (RWKV) (Peng et al., 2023), and even potentially large-scale LSTMs with sufficient memory show promise or competitive results (Schmidt, 2023). Similarly, various optimization algorithms—including Adam (Kingma and Ba, 2015), SGD, AdamW (Loshchilov and Hutter, 2019), Shampoo (Gupta et al., 2018), and the more recent Muon optimizer (Jordan et al., 2024; Lau et al., 2025)—have been effective in training these massive models. This lack of a uniquely optimal design for LLMs is reflective of the empirical trial-and-error approach driving both architectural and algorithmic innovations. Indeed, this empirical flexibility fosters co-adaptation between architectures and optimizers. Interestingly, a popular hypothesis within the AI community posits that Adam’s effectiveness may partially result from neural architectures being inadvertently “overfitted” to its optimization characteristics (Orabona, 2020).
@@ 4 Statistical Topics on LLMs
    where the reward $r(x,y)$ is trained from pairwise comparison data from human labelers using maximum likelihood estimation. The LLM is then fine-tuned using the reward model to maximize the expected reward, subject to a constraint that penalizes excessive deviation from the reference model. Owing to the stochasticity inherent in the Bradley–Terry model with noisy human feedback, this process is statistical in nature. This offers numerous opportunities for statisticians to analyze reference model misspecification, sample efficiency of preference data collection, and the generalization of learned preferences (Zhu et al., 2023; Chakraborty et al., 2024; Swamy et al., 2025; Ye et al., 2025). For example, recent studies demonstrate that the current approaches to RLHF inadequately represent the full spectrum of human preferences and can introduce statistical bias during fine-tuning (Liu et al., 2025a; Xiao et al., 2024).
@@
  • Privacy and machine unlearning: LLMs trained on vast datasets may inadvertently memorize and expose sensitive or copyrighted information. Differential privacy provides a rigorous mathematical framework that offers statistical guarantees against information leakage by introducing controlled noise during training (Dwork et al., 2006). A critical statistical consideration, therefore, is to optimize the trade-off between privacy protection and model utility, particularly for proprietary LLMs. Recent research has focused on enhancing this trade-off by implementing differential privacy during the fine-tuning phase of pre-trained models that were initially trained on public data (Li et al., 2022). However, further research is needed to meet the level of trade-off required by proprietary LLMs. A related area is machine unlearning, which aims to efficiently eliminate the influence of specific data points (e.g., due to privacy requests or copyright concerns) without retraining (Cao and Yang, 2015). This approach presents significant statistical challenges in precisely defining and verifying effective “forgetting” while preserving model capabilities (Yao et al., 2024; Zhang et al., 2024b).
@@
  • Fairness: LLMs can inherit and amplify societal biases present in their training data, leading to unfair or discriminatory outputs across different demographic groups (e.g., based on gender, race, religion) (Santurkar et al., 2023; Kotek et al., 2023). While addressing fairness is a complex socio-technical problem, statistics provides indispensable tools for defining, measuring, and mitigating bias. This includes developing quantitative fairness metrics, auditing models for systematic biases using statistical tests, and incorporating fairness considerations into various stages of the LLM pipeline—from data curation and pre-training objectives to alignment processes or even directly during generation of outputs (Zhang et al., 2024a; Chakraborty et al., 2024).
@@
  • Watermarking: To distinguish LLM-generated text from human-written content, watermarking techniques embed statistically detectable signals into the generation process based on cryptographic pseudorandomness (Kirchenbauer et al., 2023; Aaronson, 2023). Formally, the next token $w_{t+1}$ is decoded as $\mathcal{S}(\bm{P}_{t},\zeta_{t})$, where the decoder $\mathcal{S}$ is deterministic or can incorporate sampling randomness, $\bm{P}_{t}$ is the multinomial distribution for drawing the $(t+1)^{\textnormal{st}}$ token, and $\zeta_{t}$ is a pseudorandom variable that can be computed from the preceding context and a private key. The detection problem can be naturally framed as statistical hypothesis testing, based on the observation that when text is not watermarked (under the null hypothesis), $w_{t+1}$ is independent of $\mathcal{S}(\bm{P}_{t},\zeta_{t})$, while when the text is watermarked (under the alternative hypothesis), $w_{t+1}=\mathcal{S}(\bm{P}_{t},\zeta_{t})$. The latter case necessarily induces dependence between the tokens and pseudorandom variables, even without knowledge of the multinomial distributions. This framework opens avenues for applying statistical decision theory to design watermarking schemes and detection rules that achieve optimal or near-optimal detection performance while preserving text quality (Li et al., 2025). A significant practical challenge involves ensuring robustness against adversarial modifications, such as paraphrasing or translation, necessitating the development of schemes with provable statistical guarantees under such attacks (Pang et al., 2024; Li et al., 2024a). Additionally, there is substantial interest in leveraging watermarking techniques for other purposes, such as detecting data misappropriation (Cai et al., 2025).
@@
    The rationale behind this technique is that, for the larger LLM, evaluating the probability of a proposed token is computationally more efficient than decoding a token, which typically requires processing all possible tokens in the (extensively large) vocabulary. Consequently, the efficiency gain depends critically on the acceptance rate, which is inherently a statistical quantity depending on how the target LLM’s probability distributions align with those of the smaller one. Statistical analysis is crucial for optimizing the trade-off between the computational cost of the draft model and the expected speedup from the acceptance rate. For instance, this statistical technique has been implemented in DeepSeek V3 for efficient multi-token prediction (Li et al., 2024b; Liu et al., 2024) and can be integrated with watermarking (Hu and Huang, 2024). More opportunities exist for incorporating statistical insights into adaptive variations of this technique and integrating it with other methods.
@@
  • Uncertainty quantification and calibration: LLM outputs exhibit uncertainty stemming from both inherent randomness in the generation process and knowledge limitations (Yadkori et al., 2024). For trustworthy applications, particularly in high-stakes scenarios, quantifying the uncertainty of LLM outputs becomes crucial. Among various approaches addressing uncertainty, conformal prediction has emerged as a statistically rigorous and flexible methodology for providing prediction sets with distribution-free coverage guarantees—a characteristic particularly suited to the black-box nature of LLMs (Mohri and Hashimoto, 2024; Cherian et al., 2024; Gui et al., 2024). Furthermore, the utility of uncertainty estimates depends on calibration—the degree to which a model’s expressed confidence aligns with its actual accuracy. Given that aligned LLMs are often miscalibrated (Achiam et al., 2023), developing methods that simultaneously quantify uncertainty and restore calibration is an important research direction (Huang et al., 2024; Xiao et al., 2025; Wang et al., 2025; Liu et al., 2025c).
  • Evaluation: Assessing LLM capabilities across diverse tasks using benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), and GSM8K (Cobbe et al., 2021) is essential not only for tracking progress but also for guiding AI development (Silver and Sutton, 2025; Yao, 2025; Gao et al., 2025). However, the probabilistic and complex nature of LLMs introduces substantial statistical challenges in evaluation. Statistically grounded methods are required to quantify the variance and reliability of evaluation scores (Miller, 2024; Polo et al., 2024), with statistical models such as item response theory being employed for this purpose (Madaan et al., 2024). Nevertheless, this area faces an “evaluation crisis,” where reported benchmark scores frequently inflate perceived capabilities, partly due to evaluation gaming—the process of optimizing models specifically for benchmark performance (Khomenko, 2025). This phenomenon bears resemblance to statistical issues such as $p$-hacking. Consequently, we need rigorous statistical principles for robust measurement and protection against overfitting to evaluation datasets.
The central role of data.
@@
  • Data mixture and attribution: An essential challenge is determining the optimal composition of diverse data sources (e.g., web text, books, code, scientific papers) to train an LLM that achieves specific desired capabilities, often under resource constraints (Xie et al., 2023). While heuristic understanding exists—for instance, that a higher proportion of code in the training data generally leads to stronger coding abilities—the complex, high-dimensional relationship between data mixture and emergent abilities is largely unknown (Thrush et al., 2024). Statistical modeling, particularly regression-based approaches, offers a simple yet effective approach to investigating these dependencies (Liu et al., 2025b). Closely related to data mixture is the problem of data attribution, which seeks to identify the specific training samples that most influence a particular model output or behavior. Data attribution is crucial for addressing legal concerns like copyright infringement and enhancing transparency, but poses significant challenges due to the black-box nature of LLMs. Statistical techniques such as influence functions (Koh and Liang, 2017) and kernel approximation (Park et al., 2023) have been used to address these challenges for relatively small-scale models, yet significant research effort is needed to adapt and scale these methods effectively for large LLMs.
@@
    effectively captures how the pre-training loss $L$ depends on the number $N$ of model parameters and the number $D$ of training tokens, where $E$ denotes the entropy of natural text and $A,\alpha,B$, and $\beta$ are constants. These laws offer significant practical value by enabling researchers to predict potential performance gains from increased scale, thereby guiding resource allocation strategies for training progressively larger models without requiring exhaustive experimentation (Achiam et al., 2023). Scaling laws are fundamentally statistical, as they model empirically observed relationships between the scale of training resources and model performance. From a theoretical perspective, however, scaling laws present intriguing questions for statisticians. The observation that model performance continues to improve with increasing model size $N$, seemingly without saturation even when $N$ becomes extremely large, challenges classical statistical learning theories that predict model performance saturation once complexity exceeds the intrinsic dimensionality of the data. Investigating the statistical underpinnings of these empirical laws, potentially through the lens of nonparametric estimation or approximation theory in high dimensions, may yield deeper insights into the learning dynamics of LLMs and potentially lead to novel statistical methodologies.
@@
The rapidly evolving landscape of LLMs continues to generate novel statistical challenges that extend beyond the aforementioned categories, many of which remain incompletely formulated from a statistical perspective and thus present opportunities for contributions from our community. In the development of small LLMs, which are needed for deployment in edge devices, empirical evidence has revealed that knowledge distillation from larger LLMs often outperforms training from scratch (Guo et al., 2025), which calls for developing statistically efficient distillation methods. Accordingly, there is a need for owners of proprietary LLMs to develop sampling strategies that limit the effectiveness of distillation by business competitors (Savani et al., 2025). The recent emergence of reasoning models that employ latent intermediate steps via chain-of-thought (Wei et al., 2022) suggests that latent variable modeling might be valuable for understanding the mechanisms underlying the effectiveness of latent reasoning outputs (Muennighoff et al., 2025). Furthermore, the very recent advent of diffusion-based LLMs presents an opportunity for statistical analysis to elucidate the fundamental comparisons between the autoregressive and diffusion-based strategies (Nie et al., 2025; Dou et al., 2024). Moreover, as LLMs are frequently deployed as evolving API-based services, the development of statistically grounded techniques to detect unannounced updates and consequent behavioral shifts is critical for ensuring the reliability of downstream applications (Dima et al., 2025). Finally, the modification of multinomial distributions for next-token prediction through a Bayesian perspective represents another frontier where statistical insights can directly inform LLM development (Zhuang et al., 2025).

### v3

@@ 1 Introduction
Given that LLMs involve many statistical aspects, a burgeoning body of statistical research on LLMs has emerged (see a recent overview by Ji et al. (2025)). While this review paper provides a broad, tutorial-style introduction for statisticians who wish to study LLMs and obtain a big-picture view, the present paper is a shorter, more focused and selective piece that concentrates on whether LLMs would genuinely benefit from the kinds of statistical foundations that our community has developed and could further develop, rather than offering a comprehensive review. Furthermore, we ask if such statistical contributions would lead to improved development, deployment, and application, particularly to guide and enhance their real-world use? In other words, we move beyond the discussion of how LLMs can be used to enhance statistical analysis and statistical education to focus on whether statistical methodology and insights can improve LLMs themselves. [^3] For clarity, this paper does not address how LLMs can be used to enhance statistical analysis or statistical education. Interested readers are referred to Tu et al. (2024).
@@ 2 LLMs as Statistical Models
While designed as a predictive method where the underlying data-generating mechanism is treated as unknown, LLMs exhibit characteristics that distinguish them significantly from many other predictive algorithms (Breiman, 2001; Efron, 2020). Predictive methods like decision trees or support vector machines often rely on feature engineering or, equivalently, carefully chosen kernels, whereas modern LLMs operate in a nearly end-to-end data-centric manner, trained directly on vast quantities of raw data with minimal human intervention beyond data selection and cleaning. [^4] Indeed, data curation is nontrivial in practice, but from a methodological perspective, no detailed human knowledge of linguistic constructs is explicitly hard-coded into LLMs at training time. Nevertheless, human intervention remains significant regarding the choice of architectural designs for LLMs.
Owing to this strong data dependency, the capabilities of LLMs are largely determined by the properties and scale of the training data (Thrush et al., 2024). This fact is quantitatively captured by scaling laws, which demonstrate a predictable relationship between model capabilities and the volume of training data during the pre-training phase (Kaplan et al., 2020). In essence, scaling laws suggest that LLMs are what they are trained on, elevating data to the most critical component of their development. As evidence, while open-source LLM developers often release model weights and sometimes some technical details of their training processes, the composition and preparation of their training datasets are almost never disclosed (Grattafiori et al., 2024; Liu et al., 2024a).
@@
The predominant training paradigm for modern LLMs is next-token prediction (Radford et al., 2018). [^5] While other objectives exist (e.g., masked language modeling in BERT) and new approaches are emerging (e.g., diffusion models for text), the core task remains generative. Crucially, this generative process is inherently stochastic, depending not only on the prior context but also on the data used for training LLMs. This randomness is not merely an artifact but arguably a necessity for modeling human language, which itself—whether in stories, manuals, essays, creative writing, code, or mathematical proofs—is generative and rarely follows a deterministic path. This contrasts sharply with many earlier classical neural networks, particularly for classification, where outputs often correspond to fixed, deterministic ground truths.
@@
Furthermore, the variability and uncertainty stemming from the stochastic nature of LLMs demand statistical analysis. Because LLM outputs are inherently variable, understanding and quantifying the uncertainty associated with their responses is critical. This is especially true in scientific applications where reproducibility is required, and in high-stakes decision-making, such as medical diagnostics, where the model’s confidence level can drastically influence subsequent actions. Statistics offers a rich toolkit for analyzing variability, uncertainty, and miscalibration, and subsequently reducing them. Conversely, recent work demonstrated that LLM stochasticity can be harnessed to construct statistically valid inference procedures (Ratkovic, 2025).
@@ 3 LLMs as Black-box Models
The black-box nature is compounded by the fact that there is not one single “correct” architecture for LLMs achieving high performance. While the Transformer has been dominant, simplified Transformer and mixture-of-experts Transformer variants are shown empirically to work well (He and Hofmann, 2024; Dai et al., 2024). [^6] Notably, the Transformer can learn not only data patterns but also entire algorithms for statistical problems (He et al., 2025; Cao et al., 2025). Moreover, non-attention-based architectures like state-space models (Mamba) (Gu and Dao, 2024), recurrent structures (RWKV) (Peng et al., 2023), and even potentially large-scale LSTMs with sufficient memory show promise or competitive results (Schmidt, 2023). Similarly, various optimization algorithms—including Adam (Kingma and Ba, 2015), stochastic gradient descent (SGD) (Robbins and Monro, 1951), AdamW (Loshchilov and Hutter, 2019), Shampoo (Gupta et al., 2018), and the more recent Muon optimizer (Jordan et al., 2024; Lau et al., 2025; Su, 2025)—have been effective in training these massive models. This lack of a uniquely optimal design for LLMs is reflective of the empirical trial-and-error approach driving both architectural and algorithmic innovations. Indeed, this empirical flexibility fosters co-adaptation between architectures and optimizers. Interestingly, a popular hypothesis within the AI community posits that Adam’s effectiveness may partially result from neural architectures being inadvertently “overfitted” to its optimization characteristics (Orabona, 2020).
@@ 4 Statistical Topics on LLMs
    where the reward $r(x,y)$ is trained from pairwise comparison data from human labelers using maximum likelihood estimation. The LLM is then fine-tuned using the reward model to maximize the expected reward, subject to a constraint that penalizes excessive deviation from the reference model. Owing to the stochasticity inherent in the Bradley–Terry model with noisy human feedback, this process is statistical in nature. This offers numerous opportunities for statisticians to analyze reference model misspecification, sample efficiency of preference data collection, and the generalization of learned preferences (Zhu et al., 2023; Chakraborty et al., 2024; Swamy et al., 2025; Ye et al., 2025; Feng et al., 2025). For example, recent studies demonstrate that the current approaches to RLHF inadequately represent the full spectrum of human preferences and can introduce statistical bias during fine-tuning (Liu et al., 2025a; Xiao et al., 2025b). Another notable direction is to leverage experimental design for optimizing data collection for alignment, which is often costly. These approaches include adaptive data collection (Mukherjee, 2025), optimal design for reward modeling (Scheid et al., 2024), and selection of sampling strategies (Shi et al., 2025).
@@
  • Privacy and machine unlearning: LLMs trained on vast datasets may inadvertently memorize and expose sensitive or copyrighted information. Differential privacy provides a rigorous mathematical framework that offers statistical guarantees against information leakage by introducing controlled noise during training (Dwork et al., 2006). A critical statistical consideration, therefore, is to optimize the trade-off between privacy protection and model utility, particularly for proprietary LLMs. Recent research has focused on enhancing this trade-off by implementing differential privacy during the fine-tuning phase of pre-trained models that were initially trained on public data (Li et al., 2022) and generating differentially private synthetic data using LLMs without requiring model training (Xie et al., 2024). However, further research is needed to meet the level of trade-off required by proprietary LLMs. A related area is machine unlearning, which aims to efficiently eliminate the influence of specific data points (e.g., due to privacy requests or copyright concerns) without retraining (Cao and Yang, 2015). This approach presents significant statistical challenges in precisely defining and verifying effective “forgetting” while preserving model capabilities (Yao et al., 2024; Zhang et al., 2024b).
@@
  • Fairness: LLMs can inherit and amplify societal biases present in their training data, leading to discriminatory outputs across different demographic groups (Santurkar et al., 2023; Kotek et al., 2023). While addressing fairness is a complex socio-technical problem, statistics provides indispensable tools for defining, measuring, and mitigating bias. Recent empirical research has moved beyond simple disparity metrics to statistical frameworks that account for the stochastic nature of generation. For instance, Liu et al. (2024b) introduced a bias-volatility decomposition which reveals that alignment techniques like RLHF often reduce systematic bias at the cost of increasing volatility, implying that the model’s behavior becomes statistically less predictable for minority groups. Furthermore, evaluations of mitigation strategies indicate that surface-level alignment often fails to remove underlying dependencies (Bai et al., 2025). In light of this, open statistical problems remain, particularly in developing sample-efficient auditing procedures for subgroups, formulating fairness frameworks to distinguish between stereotypes and necessary semantic correlations, and incorporating fairness considerations into various stages of the LLM pipeline—from data curation and pre-training objectives to alignment processes or even directly during generation of outputs (Zhang et al., 2024a; Chakraborty et al., 2024).
@@
  • Watermarking: To distinguish LLM-generated text from human-written content, watermarking techniques embed statistically detectable signals into the generation process based on cryptographic pseudorandomness (Kirchenbauer et al., 2023; Aaronson, 2023). Formally, the next token $w_{t+1}$ is decoded as $\mathcal{S}(\bm{P}_{t},\zeta_{t})$, where the decoder $\mathcal{S}$ is deterministic or can incorporate sampling randomness, $\bm{P}_{t}$ is the multinomial distribution for drawing the $(t+1)^{\textnormal{st}}$ token, and $\zeta_{t}$ is a pseudorandom variable that can be computed from the preceding context and a private key. The detection problem can be naturally framed as statistical hypothesis testing, based on the observation that when text is not watermarked (under the null hypothesis), $w_{t+1}$ is independent of $\mathcal{S}(\bm{P}_{t},\zeta_{t})$, while when the text is watermarked (under the alternative hypothesis), $w_{t+1}=\mathcal{S}(\bm{P}_{t},\zeta_{t})$. The latter case necessarily induces dependence between the tokens and pseudorandom variables, even without knowledge of the multinomial distributions. This framework opens avenues for applying statistical decision theory to design watermarking schemes and detection rules that achieve optimal or near-optimal detection performance while preserving text quality (Li et al., 2025b). A significant practical challenge involves ensuring robustness against adversarial modifications, such as paraphrasing or translation, necessitating the development of schemes with provable statistical guarantees under such attacks (Pang et al., 2024; Li et al., 2025a). Additionally, there is substantial interest in leveraging watermarking techniques for other purposes, such as detecting data misappropriation (Cai et al., 2025).
@@
    The rationale behind this technique is that, for the larger LLM, evaluating the probability of a proposed token is computationally more efficient than decoding a token, which typically requires processing all possible tokens in the (extensively large) vocabulary. Consequently, the efficiency gain depends critically on the acceptance rate, which is inherently a statistical quantity depending on how the target LLM’s probability distributions align with those of the smaller one. Statistical analysis is crucial for optimizing the trade-off between the computational cost of the draft model and the expected speedup from the acceptance rate. For instance, this statistical technique has been implemented in DeepSeek V3 for efficient multi-token prediction (Li et al., 2024; Liu et al., 2024a) and can be integrated with watermarking (Hu and Huang, 2024). More opportunities exist for incorporating statistical insights into adaptive variations of this technique and integrating it with other methods.
@@
  • Uncertainty quantification and calibration: LLM outputs exhibit uncertainty stemming from both inherent randomness in the generation process and knowledge limitations (Yadkori et al., 2024). For trustworthy applications, particularly in high-stakes scenarios, quantifying the uncertainty of LLM outputs becomes crucial. Among various approaches addressing uncertainty, conformal prediction has emerged as a statistically rigorous and flexible methodology for providing prediction sets with distribution-free coverage guarantees—a characteristic particularly suited to the black-box nature of LLMs (Mohri and Hashimoto, 2024; Cherian et al., 2024; Gui et al., 2024). Furthermore, the utility of uncertainty estimates depends on calibration—the degree to which a model’s expressed confidence aligns with its actual accuracy. Given that aligned LLMs are often miscalibrated (Achiam et al., 2023), developing methods that simultaneously quantify uncertainty and restore calibration is an important research direction (Xiao et al., 2025a; Wang et al., 2025; Liu et al., 2025c).
  • Evaluation: Assessing LLM capabilities across diverse tasks using benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), and GSM8K (Cobbe et al., 2021) is essential not only for tracking progress but also for guiding AI development (Silver and Sutton, 2025; Yao, 2025; Gao et al., 2025). However, the probabilistic and complex nature of LLMs introduces substantial statistical challenges in evaluation. Statistically grounded methods are required to quantify the variance and reliability of evaluation scores (Miller, 2024; Polo et al., 2024; Li et al., 2025c), with statistical models such as item response theory being employed for this purpose (Madaan et al., 2024). Nevertheless, this area faces an “evaluation crisis,” where reported benchmark scores frequently inflate perceived capabilities, partly due to evaluation gaming—the process of optimizing models specifically for benchmark performance (Khomenko, 2025). This phenomenon bears resemblance to statistical issues such as $p$-hacking. Consequently, we need rigorous statistical principles for robust measurement and protection against overfitting to evaluation datasets. Techniques from experimental design can also enhance evaluation efficiency, for instance, by adaptively selecting test questions that best discriminate between model capabilities rather than using fixed benchmarks (Collins et al., 2024).
  • Hallucination: LLMs often generate fluent but factually unsupported content, a phenomenon referred to as hallucination. This is widely viewed as a “last mile” obstacle for deploying LLMs in high-stakes applications. Due to its critical importance, emerging work has been developed to detect hallucinations (Sriramanan et al., 2024; Duan et al., 2024), estimate hallucination rates (Jesson et al., 2024), and mitigate their occurrence (Huang et al., 2025). Notably, Kalai and Vempala (2024) identified a fundamentally statistical cause of hallucination: under a natural calibration condition and a Good–Turing-style approach to modeling facts, there is an inherent lower bound on the hallucination probability that does not depend on architecture or data quality (Kalai et al., 2025). While insightful, their probabilistic models remain highly stylized relative to the complexity of hallucinations observed in practice. Developing sequential probabilistic models over facts and prompts may provide a promising direction for bridging the gap between these theoretical results and the empirical behavior of hallucinations.
Central role of data.
@@
  • Data mixture and attribution: An essential challenge is determining the optimal composition of diverse data sources (e.g., web text, books, code, scientific papers) to train an LLM that achieves specific desired capabilities, often under resource constraints (Xie et al., 2023). While heuristic understanding exists—for instance, that a higher proportion of code in the training data generally leads to stronger coding abilities—the complex, high-dimensional relationship between data mixture and emergent abilities is largely unknown (Thrush et al., 2024). Statistical modeling, particularly regression-based approaches, offers a simple yet effective approach to investigating these dependencies (Liu et al., 2025b). Closely related to data mixture is the problem of data attribution, which seeks to identify the specific training samples that most influence a particular model output or behavior. Data attribution is crucial for addressing legal concerns like copyright infringement and enhancing transparency, but poses significant challenges due to the black-box nature of LLMs. Statistical techniques such as influence functions (Koh and Liang, 2017) and kernel approximation (Park et al., 2023) have been used to address these challenges for relatively small-scale models, yet significant research effort is needed to adapt and scale these methods effectively for large models (Deng et al., 2024; Pan et al., 2025).
@@
    effectively captures how the pre-training loss $L$ depends on the number $N$ of model parameters and the number $D$ of training tokens, where $E$ denotes the entropy of natural text and $A,\alpha,B$, and $\beta$ are constants. These laws offer significant practical value by enabling researchers to predict potential performance gains from increased scale, thereby guiding model design, optimal model sizing, and resource strategies for training progressively larger models without requiring exhaustive experimentation (Achiam et al., 2023). Scaling laws are fundamentally statistical, as they model empirically observed relationships between the scale of training resources and model performance. From a theoretical perspective, however, scaling laws present intriguing questions for statisticians. The observation that model performance continues to improve with increasing model size $N$, seemingly without saturation even when $N$ becomes extremely large, challenges classical statistical learning theories that predict model performance saturation once complexity exceeds the intrinsic dimensionality of the data. Investigating the statistical underpinnings of these empirical laws, potentially through the lens of nonparametric estimation or approximation theory in high dimensions, may yield deeper insights into the learning dynamics of LLMs and potentially lead to novel statistical methodologies.
Summary.
The research topics discussed above are summarized in Table 1. Broadly speaking, these directions fall into two complementary categories. The first is foundational and diagnostic, such as uncertainty quantification, hallucination, and evaluation. The second category is efficiency-driven and engineering-oriented, addressing practical needs such as data mixture optimization, watermarking, and synthetic data generation.
Table 1: Summary of statistical research directions for LLMs. The topics are grouped by their section in the text.
Topic Primary objective Statistical methodology
LLM alignment
Alignment from human feedback Preference/Control Bradley–Terry model, experimental design
Privacy & machine unlearning Security/Privacy Differential privacy, forgetting verification
Fairness Audit/Interpretability Fairness metrics
Exploiting the generative interface
Watermarking Authentication/Detection Hypothesis testing, decision theory
Speculative sampling Efficiency/Speed Rejection sampling, coupling
Tokenization Efficiency/Compression Information rate, compression algorithms
Assessment of LLM behavior
Uncertainty quantification & calibration Reliability/Diagnostics Conformal prediction, calibration
Evaluation Measurement Item response theory, experimental design
Hallucination Reliability/Factuality Good–Turing estimators, sequential models
Central role of data
Data mixture & attribution Composition/Transparency Regression, influence functions
Synthetic data & model collapse Performance/Stability Sampling theory, distributional control
Scaling laws Forecasting Nonparametric estimation, approximation theory
@@
The rapidly evolving landscape of LLMs continues to generate novel statistical challenges that extend beyond the aforementioned categories, many of which remain incompletely formulated from a statistical perspective and thus present opportunities for contributions from our community. In the development of small LLMs, which are needed for deployment in edge devices, empirical evidence has revealed that knowledge distillation from larger LLMs often outperforms training from scratch (Guo et al., 2025), which calls for developing statistically efficient distillation methods. Accordingly, there is a need for owners of proprietary LLMs to develop sampling strategies that limit the effectiveness of distillation by business competitors (Savani et al., 2025). The recent emergence of reasoning models that employ latent intermediate steps via chain-of-thought (Wei et al., 2022) offers an effective approach to improving the reasoning abilities of LLMs (Muennighoff et al., 2025). The promise of reasoning models calls for statistical methodology development in quantifying confidence for deciding when to terminate reasoning to save tokens while preserving accuracy (Sui et al., 2025; Liu et al., 2025c), as well as for performing statistical inference on the resulting reasoning outcomes, including uncertainty quantification and the assessment of internal consistency across multiple reasoning paths (Liu et al., 2025c; Wang et al., 2023). Furthermore, the very recent advent of diffusion-based LLMs presents an opportunity for statistical analysis to elucidate the fundamental comparisons between the autoregressive and diffusion-based strategies (Nie et al., 2025; Dou et al., 2024). Moreover, as LLMs are frequently deployed as evolving API-based services, the development of statistically grounded techniques to detect unannounced updates and consequent behavioral shifts is critical for ensuring the reliability of downstream applications (Dima et al., 2025). Finally, the modification of multinomial distributions for next-token prediction through a Bayesian perspective represents another frontier where statistical insights can directly inform LLM development (Zhuang et al., 2025).